Describe the use of a potentiometer for measuring
Answer #1

(A) ans

Potentiometer is used

To measure the unknown emf of the cell

To compare the emf of two cells

To determine the internal resistance of cell

To measure the current in the circuit

To measure the p.d b/w two points

To measure the thermo emf of a thermo couple

To calibrate ammeter

To calibrate voltmeter

(B) ans

A resistance can be measured by ohms law using a voltmeter and an ammeter in an electrical circuit but this measurement is not accurate to measure it more accurately

Kristie devised and Wheatstone popularized a special network design called Wheatstone bridge

It is an arrangement of four resistances which can be used to measure one of them in terms of the rest

Four resistances p Q R and S are connected as four sides of a quadrilateral ABCD as shown

A galvanometer of resistance G is connected across one set of diagonally opposite corners B and D

Condition for balancing of bridge:

Applying Kirchhoff's first law at junction B and D we get

I1-I3-Ig=0.............(1)

Applying Kirchhoff's section second law for closed loopABDA

-I1p-IgG+I2R=0,......,.......(2)

Applying Kirchhoff's section second law for closed loopBCDB

-I3Q+I4S+IgG=0..............(3)

The values of p,q,r and s are adjusted such that IG=0 at this stage the bridge is said to be in balance condition

In balanced condition the above equations respectively become

I1=I3

And I2=I4

I1P=I2R

I3Q=I4S

Dividing equation

I1P/I3Q=I2R/I4S

Using above equations we get

P/Q=R/S

This is the balancing condition for Wheatstone's bridge

The functions of Wheatstone's bridge

1 we can compare two unknown resistances R and S from

P/Q=R/S

2 if the value of S is also known we can calculate the value of unknown resistance R

3 the small strains produced in hard materials can be measured accurately with a Wheatstone bridge using a strain guage

(C) ans

The process of raising the strength of a weak input signal to strong output signal is called amplification

the circuit of amplifier

Amplifier :

This device which increases the weak input signal into strong output signal is called an amplifier

The function of amplifier

The amplifier is mostly used in common emitter circuit

Amplifier is used to find the voltage of output signal

Amplifier is used in fire alarms

Amplifier is used in medical industry

Amplifier is used in to find depth of sea